Minutes: November 3 2011
Attendence:Carol, Kimberly, Arielle, Dong-Chan, Mike, Aaron, Scott, Jonathan, Tony, Lyndele,
1. News and Main Items
- Data Management Review - overview from Mark Lacy
- Socorro, GB and NAASC are undergoing a data management review - NAASC has developed a data management plan. Plan covers everything "once data hits the disc."
2. Task leads reports / new business
- Site, CSV (Al, staff in Chile)
- HIA (Gerald)
- received ideas for Primer; postdocs in ESO will put together new science cases that use the new Cycle 1 capabilities.
- Action (Gerald): should be in contact with ESO postdocs to make sure that cases are coordinated.
- Attended the P2G bootcamp
- actively working batch 2 project; begun on a batch 3 project. CS training for staff
- as of April 1, will no longer be called the Hertzburg Institute - not yet known what they will be called.
- Taiwan & EA Liaison (Aaron)
- SD school will be Dec 7,8. Agenda to be similar to the one used at EU ARC. Noted that this was high level and we'd like to have SD data to work with
- Documentation (Jim, Jonathan)
- Action from last week: Jim B to contact Eelco about science examples for the Primer
- Jonathan: Requesting great mm astrophysics (text and/or slides) that Jonathan and Jim can use as a basis for course, rather than starting from scratch.
- Focus: perhaps more useful to develop a course that is aimed at a particular audience, and hold it as a summer school/graduate level course? submm/mm astrophysics & observing techniques.
- Adam - suggested IRAM model for summer courses.
- Helpdesk (Tony)
- Action from last week: Please use "Pending" if still working with PI instead of "On hold"
- PI Support & Visitors (Scott, Carol)
- Scott: CSs can ask PIs to approve SBs pending finding a calibrator; AoDs will find the calibrators. Only one or two batch 2 cases where SBs have not been generated.
- NOTE post meeting: may need to have all CSs carefully review all calibrators - more information to follow
- Scott will review the tickets to make sure that calibrater issue is not holding things up.
- Q on how f2f visits will be arranged; answer: helpdesk
- The webform survey for proposers nearing completion - led by ESO
- Workshops, Schools, Meetings (Tony, Adam)
- Dec 1-2 data processing workshop
- Action from last week: slides due ahead for review & dry run on Nov 15
- Summer school
- Outflows, Winds and Jets
- Oct 8-12 ALMA Works - "first results" in Santiago.
- Tutorials (Carol)
- Community Outreach and Science Web (Mark, Aaron)
- Action from last week: Aaron is planning to write a 'policy' document that describes the purpose/process of the website. Should be ready by next USG meeting. Dong-Chan is getting access to Tania's pictures from the site. Once EPO page is up and running, we'll simply retrieve pages from that site.
- Computing (Mike)
- database synchronization is about 1 week out. Some Oracle problems for EA and NA; EU is OK
- ESO tool for extracting data from archive; we did some work on this at NAASC. Still issue with some things missing in the archive
- next phase is integration into web access
- DSG liaisons (Adam, Tony)
- JSG liaison (John)
3. Current Focus Items
- Batch 2 SB preparation time period ended; P2G members proceed to Batch 3, though Jan 10.
- Single Dish training
- Cycle 1 capabilities, configurations, requirements on tools, simulations, documentation, staff training
- CDEs: materials and dry runs
- 1-2 December Data Processing Workshop: materials and dry run Nov 15
- AAS dry runs, Dec 12